OpenStudy (anonymous):

Expand and simplify.(i reprsents the imaginary unit ) (i-3)^4

OpenStudy (anonymous):

(i-3)^4 =(i-3)^2 (i-3)^2 =(i^2-6i+9)(i^2-6i+9) where i^2=-1 =(-1-6i+9)(-1-6i+9) =(8-6i)(8-6i) =(8-6i)^2 =(64-48i-48i+36i^2) =64-96i-36 =28-96i ans....